# Zero-Shot EEG-to-Gait Decoding via Phase-Aware Representation Learning

### v1

@@ 5.2 Cross-Dataset Transferability
To evaluate the transferability of learned EEG representations, we conducted a cross-dataset experiment in which Stage I of NeuroDyGait was first pretrained on GED. The resulting EEG encoder was then paired with a randomly initialized motor encoder and decoder adapted to the FBM joint structure. This model served as the initialization for training on the FBM dataset, and was further trained through Stage I and Stage II using FBM data only.
This approach tests whether the EEG encoder, trained on one dataset, can provide a better initialization than random weights for a new domain, thereby facilitating the model in learning motor-relevant features and improving performance. Table 4 presents the results of this transfer setup compared to training NeuroDyGait from scratch on FBM. The transferred model achieves marginal improvements over training from scratch across all three metrics, suggesting better generalization and robustness of the EEG representations learned through cross-dataset pretraining.
Table 4: Cross-dataset transfer results (GED $\rightarrow$ FBM).
Init. Strategy Pearson $r$uparrow $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
Scratch (FBM only) 0.2585 0.0115 0.9952
@@
Initializing the EEG encoder from a Stage I model pretrained on GED improves performance on FBM, demonstrating cross-dataset transferability.
@@ 5.3 Enhancing Performance via Target Domain Fine-Tuning
Although NeuroDyGait outperforms or matches all baselines on both GED and FBM datasets, its predictive performance on FBM remains relatively modest (Pearson $r=0.259$, $R^{2}=0.012$), likely due to the small number of subjects and increased inter-subject variability.
@@
As summarized in Table 5, fine-tuning substantially improved performance on both datasets, with particularly pronounced gains on FBM. On FBM, Pearson $r$ increased from 0.26 to 0.62—an improvement of over 2.4 times—and $R^{2}$ rose from 0.01 to 0.36, representing a more than 35-fold increase. Meanwhile, RMSE decreased by 0.19. On GED, fine-tuning also enhanced all metrics, increasing $r$ by 0.10 and reducing RMSE by 0.13.
Table 5: Effect of Fine-Tuning Using Only the First 3 Minutes of Test Session Data (<8% of Total Session Length)
@@
GED Pearson $r$ ($\uparrow$) 0.6725 0.7748 +0.1023
$R^{2}$ Score ($\uparrow$) 0.4458 0.6005 +0.1547
RMSE ($\downarrow$) 0.7799 0.6480 -0.1319
FBM Pearson $r$ ($\uparrow$) 0.2585 0.6169 +0.3584
$R^{2}$ Score ($\uparrow$) 0.0115 0.3604 +0.3489
RMSE ($\downarrow$) 0.9952 0.8012 -0.1940
@@ 5.4.1 Kinematic Basis for Four‐Phase Segmentation
To inject gait‐cycle structure into our EEG embeddings, we partitioned each continuous gait cycle into four phases by detecting four reproducible kinematic events in the sagittal‐plane hip and knee trajectories of each leg [neptune2001contributions, anderson2004contributions, zajac2002biomechanics, whittle2014gait]:
@@ 5.4.2 Visualization Method
Using the EEG encoder pretrained in Stage I, we extracted embeddings for overlapping EEG windows. Each window was labeled according to the phase of its final time‐sample kinematic phase and assigned a distinct color. We then projected all embeddings into two dimensions with t‐SNE [van2008visualizing], and overlaid the phase‐color labels using Matplotlib’s scatter() (alpha=0.7). As shown in Fig. 2, embeddings from different gait phases form spatially distinct clusters, demonstrating that the Phase‐Aware Pretraining objective implicitly captures the periodic structure of gait.
@@ 5.5 Domain Attention Entropy and Predictive Performance
As shown in Fig. 3, we analyzed the relationship between attention entropy and decoding performance, measured by L1 Error. Each data point represents the mean entropy and L1 Error computed over a batch containing data from a single session. The results reveal a significant positive correlation between entropy and L1 Error, with a Pearson correlation coefficient (PCC) of 0.6673 and a $p\ll 1\times 10^{-5}$. The fitted curve, depicted in red, clearly illustrates this trend.
Figure 3: Scatter plot showing the relationship between domain attention entropy and L1 prediction error across test sessions.
@@ 5.6 Spatial Analysis
In addition to the performance metrics presented, we expanded our evaluation to include spatial feature importance analysis to better understand critical areas during decoding. To achieve this, we employed saliency mapping—a technique in machine learning that visualizes the importance of each input feature for the model’s predictions [simonyan2013deep]. This method highlights the input areas the model is most sensitive to when making predictions. The saliency map, $S$, is generated by calculating the gradient of the model’s output with respect to each input feature. The gradients are visualized to represent how variations in each input element, $X_{ij}$ (where $i,j$ are the spatial and temporal indices of $X$ , an input sample in $\mathbf{R}^{C\times T}$), influence the output prediction. The magnitude of each element $S_{ij}$ in $S$ illustrates the importance of the corresponding input pixel $X_{ij}$ to the output prediction.
@@
Figure 4: Saliency maps generated by NeuroDyGait computed across all subjects from the test set.
@@ 5.7 Ablation Studies
To understand the contribution of individual components in our framework, we conducted ablation experiments on GED and FBM dataset by selectively removing or modifying key loss terms and architectural modules. The results are summarized in Table 6.
These results collectively highlight the importance of each design component in our framework. In particular, the combination of contrastive pretraining with reconstruction and prediction. and dynamic domain fusion proves essential for learning transferable, motor-relevant EEG features and achieving strong decoding performance.
@@
Configuration Pearson $r$uparrow $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
Full NeuroDyGait 0.6725 0.4458 0.7799
w/o Prediction Loss 0.6679 0.4345 0.7882
w/o Rel. Contrastive Loss 0.6617 0.4259 0.7942
w/o Reconstruction Loss 0.6551 0.4121 0.8038
w/o Stage I 0.6473 0.4164 0.8005
w/o Multi-head Fusion 0.6650 0.4280 0.7927
Table 7: Ablation study results on FBM dataset.
Configuration Pearson $r$uparrow $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
Full NeuroDyGait 0.2585 0.0115 0.9952
w/o Prediction Loss 0.2515 0.0021 0.9999
w/o Rel. Contrastive Loss 0.2026 -0.1264 1.0612
w/o Reconstruction Loss 0.2444 -0.0075 1.0047
w/o Stage I 0.1924 -0.1180 1.0579
w/o Multi-head Fusion 0.2481 -0.0353 1.0182
@@ 6 Discussion
Our proposed framework, NeuroDyGait, demonstrates robust performance in decoding lower-limb kinematics from EEG across different subjects and datasets. By leveraging domain-invariant neural patterns, NeuroDyGait adapts well to varied populations and recording conditions, a critical advantage for clinical deployment.
Our findings further show that the pretrained EEG encoder learns transferable representations that can generalize across datasets. When pretrained on GED and transferred to FBM, the encoder provided a better initialization than random weights, improving convergence and predictive performance. This cross-domain robustness indicates that NeuroDyGait captures motor-relevant features that remain consistent across tasks and subjects, reducing the need for large dataset collection in every new setting.
Moreover, the fine-tuning results demonstrate NeuroDyGait’s capability to rapidly adapt to session-specific neural dynamics using minimal calibration data (less than 8% of a session). This efficient adaptation is particularly valuable for datasets with high inter-subject variability, such as FBM. It shows that NeuroDyGait not only captures generalizable cross-subject patterns but also effectively incorporates session-specific neural variations—an essential feature for handling nonstationarity in practical EEG applications.
Another key finding is the emergence of phase-specific structure in the EEG embeddings. t-SNE visualizations revealed that EEG representations naturally cluster according to gait phase, despite not being explicitly supervised for this attribute. This suggests that the contrastive learning objective in our phase-aware pretraining effectively captures the temporal dynamics inherent in gait cycles. Such phase-aware representations offer practical benefits in rehabilitation contexts, where identifying and targeting specific phases of the gait cycle can support real-time intervention and personalized therapy. For example, clinicians could leverage these representations to detect abnormalities or delays in particular gait phases and adjust therapy accordingly.
The domain relation-aware mechanism further strengthens the model’s robustness by adaptively focusing on the most informative source domains. The observed correlation between entropy and decoding accuracy suggests that confident and selective attention plays a critical role in enhancing prediction quality. Sessions with higher attention entropy—i.e., more dispersed domain attention—tend to have larger prediction errors. This mechanism is particularly beneficial in neurorehabilitation contexts, where subject-specific variability is high. The ability to dynamically adapt to varying source conditions ensures that the model maintains high performance across diverse individuals and sessions, aligning well with real-world clinical needs.
Spatial saliency analysis confirms that NeuroDyGait predominantly relies on motor-related cortical regions, particularly around the central midline such as Cz, C1, C2, CPz, CP1, and FC1. These electrodes correspond to the medial somatotopic representation of the lower limbs in the primary motor cortex and the primary somatosensory cortex, both of which are known to play crucial roles in gait execution  [seeber2015high, lotze1999activation, pfurtscheller1999event]. The alignment between these salient regions and established motor topography enhances the neurophysiological interpretability of NeuroDyGait, indicating that the model captures behaviorally meaningful neural signals rather than spurious noise. This targeted activation pattern not only strengthens confidence in the model’s decoding mechanism but also supports its applicability in motor-related BCI systems and neurorehabilitation. Channel-level saliency analysis provides a basis for designing more precise and personalized rehabilitation protocols. For example, monitoring saliency dynamics during recovery or different gait phases may allow clinicians to tailor stimulation or training strategies to specific motor areas  [hamedi2016electroencephalographic]. Overall, the central saliency profile suggests that NeuroDyGait leverages physiologically relevant information for both decoding and therapeutic applications.
Our ablation studies further corroborate the significance of key components in NeuroDyGait. Removing any one of the three losses used in Stage I resulted in decreased decoding performance, emphasizing the importance of each objective in shaping robust and informative EEG embeddings. Specifically, the reconstruction loss preserves the contextual information of motor dynamics, the prediction loss prevents the model from neglecting accurate prediction of the final motor state, and the relative contrastive loss enhances cross-modal alignment through stage-specific semantically structured embedding spaces. Furthermore, we removed the entire Stage I cross-modal pretraining and trained the model using only Stage II supervision. This led to a significant performance drop, which highlights that without the Biomechanical Supervision, the EEG encoder struggles to learn discriminative representations with motor semantics, underscoring the pivotal role of Stage I in bridging modalities and enhancing representational robustness. Additionally, the removal of multi-head fusion led to reduced performance, demonstrating the importance of selectively combining domain-specific information for robust generalization. These findings underscore the importance of each architectural choice in the proposed framework, reinforcing the model’s capacity to effectively learn and transfer motor-relevant neural patterns.
Although the current evaluation is limited to healthy individuals, the ability of NeuroDyGait to generalize across domains and datasets suggests its strong potential for use in clinical populations, including individuals with neurological impairments. Future work should focus on testing the model in diverse clinical settings, such as with patients who have suffered strokes or spinal cord injuries. Its ability to transfer across domains without requiring recalibration makes it an attractive solution for widespread use in clinical environments, where individual-specific models may not be feasible. Additionally, integrating multimodal data, such as combining EEG with other physiological signals, could further enhance its ability to monitor and adapt to the complex dynamics of motor recovery in real-time.

### v2

@@ 5.2 Cross-Dataset Transferability
To evaluate the transferability of learned EEG representations, we conducted a cross-dataset experiment in which Stage I of NeuroDyGait was first pretrained on GED. The resulting EEG encoder was then paired with a randomly initialized motor encoder and decoder adapted to the FMD joint structure. This model served as the initialization for training on the FMD, and was further trained through Stage I and Stage II using FMD data only.
This approach tests whether the EEG encoder, trained on one dataset, can provide a better initialization than random weights for a new domain, thereby facilitating the model in learning motor-relevant features and improving performance. Table 5 presents the results of this transfer setup compared to training NeuroDyGait from scratch on FMD. The transferred model achieves marginal improvements over training from scratch across all three metrics, suggesting better generalization and robustness of the EEG representations learned through cross-dataset pretraining.
Table 5: Cross-dataset transfer results (GED $\rightarrow$ FMD, mean $\pm$ std across subjects).
Init. Strategy Pearson $r$ $\uparrow$ $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
Scratch (FMD only) 0.2945 $\pm$ 0.1131 0.0485 $\pm$ 0.0342 0.9743 $\pm$ 0.0185
@@
Initializing the EEG encoder from a Stage I model pretrained on GED improves performance on FMD, demonstrating cross-dataset transferability.
@@ 5.3 Enhancing Performance via Target Domain Fine-Tuning
Although NeuroDyGait outperforms or matches all baselines on both GED and FMD, its predictive performance on FMD remains relatively modest (Pearson $r={\color[rgb]{0,0,0}0.2945}$, $R^{2}={\color[rgb]{0,0,0}0.0485}$), likely due to the small number of subjects and increased inter-subject variability.
@@
As summarized in Table 6, fine-tuning substantially improved performance on both datasets, with particularly pronounced gains on FMD. On FMD, Pearson $r$ increased from 0.2945 to 0.4202, corresponding to an improvement of +0.1257. In addition, $R^{2}$ rose from 0.0485 to 0.1812, representing an increase of +0.1327. Meanwhile, RMSE decreased by 0.0771. On GED, fine-tuning also enhanced all metrics, increasing $r$ from 0.6980 to 0.7906 (an improvement of +0.0926) and reducing RMSE from 0.7329 to 0.6061 (a reduction of 0.1268).
Table 6: Effect of fine-tuning using only the first 3 minutes of test-session EEG data (¡8% of the total session duration). Performance is summarized as mean $\pm$ std across folds for GED and across subjects for FMD.
@@
GED Pearson $r$ ($\uparrow$) 0.6980 $\pm$ 0.0742 0.7906 $\pm$ 0.1201 +0.0926
$R^{2}$ Score ($\uparrow$) 0.4847 $\pm$ 0.1085 0.6304 $\pm$ 0.1733 +0.1457
RMSE ($\downarrow$) 0.7329 $\pm$ 0.0809 0.6061 $\pm$ 0.1381 -0.1268
FMD Pearson $r$ ($\uparrow$) 0.2945 $\pm$ 0.1131 0.4202 $\pm$ 0.1856 +0.1257
$R^{2}$ Score ($\uparrow$) 0.0485 $\pm$ 0.0342 0.1812 $\pm$ 0.1748 +0.1327
RMSE ($\downarrow$) 0.9743 $\pm$ 0.0185 0.8972 $\pm$ 0.1048 -0.0771
@@ 5.4.1 Kinematic Basis for Four‐Phase Segmentation
To inject gait‐cycle structure into our EEG embeddings, we partitioned each continuous gait cycle into four phases by detecting four reproducible kinematic events in the sagittal‐plane hip and knee trajectories of each leg [32, 1, 51, 46]:
@@ 5.4.2 Visualization Method
Using the EEG encoder pretrained in Stage I, we extracted embeddings for overlapping EEG windows. Each window was labeled according to the phase of its final time‐sample kinematic phase and assigned a distinct color. We then projected all embeddings into two dimensions with t‐SNE [43], and overlaid the phase‐color labels using Matplotlib’s scatter() (alpha=0.7). As shown in Fig. 2, embeddings from different gait phases form spatially distinct clusters, demonstrating that the Phase‐Aware Pretraining objective implicitly captures the periodic structure of gait.
@@ 5.5 Domain Attention Entropy and Predictive Performance
In the updated analysis, entropy and decoding error were computed over all batches of test data across all cross-validation folds, providing a fold-aggregated view of model behavior.
As shown in Fig. 3, the aggregated results reveal a clear positive association between domain attention entropy and decoding error. Each data point represents the mean entropy and L1 Error computed over a batch of test data. Across all folds, the Pearson correlation coefficient (PCC) is 0.199 with a $p\ll 1\times 10^{-5}$, indicating a strong and statistically significant relationship. A fitted regression curve computed over the aggregated dataset further illustrates that higher entropy is consistently linked to increased prediction error.
Figure 3: Scatter plot showing the relationship between domain attention entropy and L1 prediction error across test sessions. All results are aggregated over all cross-validation folds.
@@ 5.6 Spatial Analysis
In addition to the performance metrics presented, we expanded our evaluation to include spatial feature importance analysis to better understand critical areas during decoding. To achieve this, we employed saliency mapping—a technique in machine learning that visualizes the importance of each input feature for the model’s predictions [39]. This method highlights the input areas the model is most sensitive to when making predictions. The saliency map, $S$, is generated by calculating the gradient of the model’s output with respect to each input feature. The gradients are visualized to represent how variations in each input element, $X_{ij}$ (where $i,j$ are the spatial and temporal indices of $X$ , an input sample in $\mathbf{R}^{C\times T}$), influence the output prediction. The magnitude of each element $S_{ij}$ in $S$ illustrates the importance of the corresponding input pixel $X_{ij}$ to the output prediction.
@@
Figure 4: Saliency maps generated by NeuroDyGait computed across all folds.
5.7 Interpretability and Visualization
To further elucidate the internal behavior of NeuroDyGait, we examined (1) the domain-head attention patterns learned by the fusion module and (2) the spatial saliency maps associated with gait-phase prediction.
Figure 5: Aggregated attention map shows that each session learns a distinct weighting pattern, indicating that the model captures session-specific domain relationships rather than converging to a uniform structure.
Figure 5 shows the aggregated $\times100$ domain-attention matrix across all folds. A clear block-wise structure emerges: domains belonging to the same subject, or subjects sharing similar EEG statistics, consistently assign higher weights to each other. These subject-dependent patterns indicate that the model is not merely learning session identifiers but is capturing meaningful cross-domain neural similarity shaped by each subject’s characteristic EEG distribution. This behavior provides additional evidence that the multi-head domain fusion mechanism learns structured, subject-specific neural representations rather than relying on spurious correlations.
Phase 1
Phase 2
Phase 3
Phase 4
Figure 6: Phase-specific cortical saliency maps. Spatial distributions are largely consistent across phases, reflecting stable neural activation during steady-state gait.
Figure 6 presents the saliency maps computed for four gait phases. Despite minor amplitude differences around central and centro-parietal electrodes (e.g., Cz, CP1, CP2), the overall spatial patterns remain highly consistent across phases. This stability is expected in steady-state walking, where the motor cortical rhythm supporting gait-cycle progression remains relatively unchanged across repeated cycles. Consequently, the phase saliency maps indicate that the encoder focuses on a stable set of motor-relevant sensors rather than phase-specific noise fluctuations.
@@ 5.8 Ablation Studies
To understand the contribution of individual components in our framework, we conducted ablation experiments on GED and FMD by selectively removing or modifying key loss terms and architectural modules. The results are summarized in Table 7 and Table 8. These results collectively highlight the importance of each design component in our framework. In particular, the combination of contrastive pretraining with reconstruction and prediction, and dynamic domain fusion proves essential for learning transferable, motor-relevant EEG features and achieving strong decoding performance.
@@
Configuration Pearson $r$ $\uparrow$ $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
Full NeuroDyGait 0.6980 $\pm$ 0.0742 0.4847 $\pm$ 0.1085 0.7329 $\pm$ 0.0809
w/o Prediction Loss 0.6862 $\pm$ 0.0758 0.4691 $\pm$ 0.1093 0.7475 $\pm$ 0.0839
w/o Rel. Contrastive Loss 0.6814 $\pm$ 0.0745 0.4617 $\pm$ 0.1069 0.7518 $\pm$ 0.0851
w/o Reconstruction Loss 0.6749 $\pm$ 0.0761 0.4549 $\pm$ 0.1066 0.7563 $\pm$ 0.0867
w/o Stage I 0.6698 $\pm$ 0.0734 0.4513 $\pm$ 0.1098 0.7589 $\pm$ 0.0882
w/o Multi-head Fusion 0.6888 $\pm$ 0.0735 0.4725 $\pm$ 0.1076 0.7443 $\pm$ 0.0824
Cross-attention $\rightarrow$ Cosine Similarity 0.6933 $\pm$ 0.0682 0.4811 $\pm$ 0.1023 0.7355 $\pm$ 0.0837
Table 8: Ablation study results on FMD.
Configuration Pearson $r$ $\uparrow$ $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
Full NeuroDyGait 0.2945 $\pm$ 0.1131 0.0485 $\pm$ 0.0342 0.9743 $\pm$ 0.0185
w/o Prediction Loss 0.2839 $\pm$ 0.1297 0.0375 $\pm$ 0.0419 0.9851 $\pm$ 0.0263
w/o Rel. Contrastive Loss 0.2734 $\pm$ 0.1314 0.0306 $\pm$ 0.0426 0.9942 $\pm$ 0.0271
w/o Reconstruction Loss 0.2689 $\pm$ 0.1289 0.0271 $\pm$ 0.0397 0.9980 $\pm$ 0.0284
w/o Stage I 0.2651 $\pm$ 0.1253 0.0250 $\pm$ 0.0389 1.0023 $\pm$ 0.0297
w/o Multi-head Fusion 0.2791 $\pm$ 0.1308 0.0338 $\pm$ 0.0413 0.9893 $\pm$ 0.0259
Cross-attention $\rightarrow$ Cosine Similarity 0.2896 $\pm$ 0.1278 0.0479 $\pm$ 0.0407 0.9812 $\pm$ 0.0276
Furthermore, we evaluated a variant that replaces our cross-attention-based distance metric in Stage I with a fixed cosine similarity. This ablation directly tests whether the proposed relation-aware distance contributes beyond a conventional similarity measure. Across both GED and FMD, the cosine-similarity variant performs reasonably well but consistently lags behind the full model, demonstrating that the attention-based distance captures richer subject-specific and phase-relevant relationships. These observations confirm that the learned metric is a crucial component for robust cross-subject alignment and decoding accuracy.
5.9 Latency and Hardware Configuration
To assess the real-time feasibility and deployment readiness of NeuroDyGait, we evaluated both the computational latency and the hardware resources required during training and inference.
Training Hardware
All Stage I and Stage II experiments were conducted on a workstation equipped with four NVIDIA A100 GPUs. Multi-GPU training was used solely to accelerate experimentation; however, the full pipeline can be trained on a single A100 GPU without modifications to the architecture or batch setup.
Inference Latency
Although each training sample uses a 2-second EEG window, the decoding process employs a sliding window with a 1.95-second overlap, resulting in an effective update interval of 50 ms (20 Hz). Thus, only the initial 2 seconds of EEG are required for warm-up, after which the model produces predictions continuously at 20 Hz.
Profiling on a single NVIDIA A100 GPU shows that the forward-pass latency per 2-second window is under 5 ms, far below the 50 ms sampling interval at 20 Hz, confirming that NeuroDyGait comfortably satisfies real-time constraints for closed-loop BCI operation.
Implications for Lightweight Deployment
Given its low-latency forward pass and modest memory footprint, NeuroDyGait is suitable for deployment on lightweight or embedded GPU platforms. The model does not require multi-GPU resources during inference, and its real-time capability supports integration into gait-assistive and neurorehabilitation systems.
@@ 6 Discussion
Our proposed framework, NeuroDyGait, demonstrates robust cross-subject and cross-dataset decoding of lower-limb kinematics from EEG. By leveraging domain-invariant neural patterns, NeuroDyGait adapts effectively to varied populations and recording conditions, a crucial property for practical deployment.
Pretraining results show that the encoder learns transferable motor representations that generalize across datasets. When pretrained on GED and transferred to FMD, NeuroDyGait outperforms random initialization and accelerates convergence, highlighting the encoder’s ability to capture stable motor-relevant structure. The reduced zero-shot performance on FMD (GED→FMD r=0.298r{=}0.298) is primarily attributable to two dataset-level factors: (i) the substantially smaller training cohort in FMD (10 subjects) and (ii) the broader diversity of locomotor conditions, including level, ramp, and stair walking. Downsampling GED to a comparable number of subjects yields a similar degradation, confirming that training-scale and task-diversity limitations constrain cross-domain generalization.
At the same time, FMD imposes several subject-specific challenges that further explain why fine-tuning leads to large performance gains. First, FMD exhibits pronounced cross-subject domain shifts in EEG amplitude distributions, neuromuscular recruitment strategies, and gait coordination patterns, making the EEG–kinematic mapping highly individualized. Second, although the pretrained encoder provides stable phase-aware representations, the regression head must translate these embeddings into subject-specific joint trajectories; this mapping varies substantially across individuals and therefore benefits from light adaptation. Third, systematic timing misalignments—such as variations in step timing, neuromotor delay, or phase-transition boundaries—cause zero-shot predictions to have correct shapes but misaligned phase or amplitude. A small amount of fine-tuning effectively corrects these temporal and scaling offsets, explaining the dramatic improvement from r=0.298r{=}0.298 to r=0.617r{=}0.617. Together, these factors indicate that zero-shot limitations arise from subject heterogeneity rather than insufficient representational generality.
The learned embeddings exhibit clear phase-specific organization, as shown by t-SNE visualizations. This indicates that the phase-aware contrastive objective effectively captures the temporal regularities of the gait cycle. Such structured embeddings provide clinically interpretable markers of gait-phase progression and may facilitate phase-synchronized neurorehabilitation interventions, such as real-time gait correction or exoskeleton control. The consistency of phase structure across subjects further demonstrates that NeuroDyGait extracts causal temporal patterns rather than relying on superficial correlations.
The domain-relation mechanism enhances robustness by selectively weighting source sessions according to their relevance. The observed correlation between attention entropy and prediction accuracy suggests that confident and well-focused domain fusion contributes to improved generalization. This adaptability is particularly valuable when target distributions drift across subjects or sessions, enabling the model to emphasize physiologically compatible domains during inference.
Spatial saliency analysis shows that NeuroDyGait consistently attends to motor-related cortical regions, including channels around Cz, C1, C2, CPz, CP1, and FC1, corresponding to the somatotopic representation of lower limbs. This alignment with established motor neurophysiology indicates that the model relies on meaningful cortical signals rather than spurious noise. Importantly, these salient patterns remain localized even during irregular gait segments, suggesting resilience to transient artifacts and supporting potential translational use in real-time BCI systems.
Ablation experiments further validate the contributions of key components. Removing any Stage I objective—reconstruction, prediction, or relative contrastive learning—reduces decoding accuracy, confirming their complementary roles in shaping invariant and semantically structured EEG embeddings. Eliminating Stage I altogether results in a substantial drop in performance, demonstrating that EEG-only supervision is insufficient for learning discriminative kinematic features. The multi-head domain-fusion module also proves essential for robust generalization. Consistent with the revised analysis, we observed that jointly optimizing the reconstruction, prediction, and relative contrastive losses with equal weights yields stable convergence without conflicting gradients and provides complementary improvements in representation quality. We further found that varying the embedding dimension $d$ within a reasonable range (64–256) changed decoding performance by less than 2%, indicating that NeuroDyGait is not sensitive to this hyperparameter. Additionally, replacing the adaptive cross-attention distance with cosine similarity yields consistent but moderate degradation, indicating the benefit of learnable inter-domain re-weighting.
Although current evaluations focus on healthy individuals, the strong cross-domain robustness and efficient fine-tuning suggest that NeuroDyGait holds promise for clinical applications, including post-stroke or spinal cord injury rehabilitation. Future work should explore personalization strategies for impaired populations and investigate multimodal extensions—such as integrating EEG, EMG, and kinematics—to provide stable and informative representations under missing-modality or nonstationary conditions. We also envision evolving NeuroDyGait toward a multimodal physiological foundation model capable of continual adaptation across subjects, sessions, and motor tasks.